OpenStudy (anonymous):

can anybody please explain the steps of solving this equation for me? n/5 - 3n/10 = 1/5

OpenStudy (saifoo.khan):

solve the lleft side by taking LCM

OpenStudy (saifoo.khan):

left*

OpenStudy (anonymous):

2n/10 - 3n/10 =1/5 (2n-3n)/10 =1/5 = 2/10 both sides *10 : 2n-3n=2 so -n=2 and n=-2